OpenStudy (anonymous):

Use the function to evaluate the indicated expressions and simplify. f(x) = -8x^2 + 2 f(x + 2) = _____ f(x) + f(2) = _____

OpenStudy (lalaly):

f(x+2) = -8(x+2)^2 +2= -8[x^2+4x+4]+2=-8x^2 -32x-30 f(x)+f(2)= -8x^2 +2 -32+2 =-8x^2-28